Solve for x:
8(^2x)+8^x-12=0

Answer by edjones(8007)

8^(2x)+8^x-12=0
(8^x)^2+8^x-12=0
(8^x+4)(8^x-3)=0 Factor
.
8^x-3=0
8^x=3
x=log[8]3
x=.528321
.
Let me know if you don't understand any of this.
.
Ed